OpenStudy (anonymous):

E and F are events in a sample space S for which P(E) = 0.3, P(F) = 0.5, P(E F) = 0.7 Find the probability Round your answer to 2 decimal places.

OpenStudy (anonymous):

Find the probability of what P(E/F) of P(F/E)?

OpenStudy (anonymous):

P(F)

OpenStudy (anonymous):

I'm confused.......is this a conditional probability problem? because if you looking for P(F) you already have it, but if your looking for the probability of P(F) give the probability of P(E) you have to use bayes formula

OpenStudy (anonymous):

yes its a conditional probablity problem

OpenStudy (anonymous):

thought so...... its going to be bayes formula P(F/E)=P(FE)/P(E) P(F/E)=0.7/0.3 because P(FE)=P(EF)